OpenStudy (anonymous):

x^3-12x^2+48x-64=0 pls help

OpenStudy (anonymous):

The only solution is x=4.

OpenStudy (anonymous):

Because the left side equals (x-4)^3.

OpenStudy (anonymous):

show step pls

OpenStudy (anonymous):

well, its simple because x^3-12x^2+48x-64=(x-4)*(x-4)*(x-4) = (x-4)^3.

OpenStudy (anonymous):

You can see this by factoring out (x-4)^3. Or you can perform polynomial division of your polynomial by x-4. And then divide the result by x-4 again which should give you x-4 as the result, proving that the polynomial equals (x-4)^3.

OpenStudy (anonymous):

ok i see.thanks.